Solve (sinx)sec^2(x)+(cos x)cosec^2 (x)=0.


1
Expert's answer
2022-02-01T10:27:55-0500

Calculation.


"\qquad\qquad
\begin{aligned}
\small \sin\theta.\sec^2\theta+\cos\theta.\csc^2\theta&=\small 0\\
\small \frac{\sin\theta}{\cos^2\theta}+\frac{\cos\theta}{\sin^2\theta}&=\small 0\\
\small \frac{\sin^3\theta+\cos^3\theta}{\sin^2\theta.\cos^2\theta}&=\small0\\
\\
\small \sin^2\theta.\cos^2\theta&\not=\small 0\\
\\
\therefore\,\small \sin^3\theta+\cos^3\theta&=\small 0\\
\small (\sin\theta+\cos\theta)(1-\sin\theta.\cos\theta)&=\small 0\\
\\
\small 1-\sin\theta.\cos\theta&=\small 0\\
\small \sin2\theta&=\small 2\\
\because \,\small -1\leq&\small \sin2\theta\leq1\\
\therefore\,\text{neglected},\\
\\
\small \sin\theta+\cos\theta&=\small 0\\
\small \tan\theta&=\small -1\cdots(\cos\theta\not=0)\\
\small \theta&=\small \frac{3\pi}{4}+n\pi\cdots\cdots(\text{Answer})\\

\end{aligned}"